The first and the last term of an arithmetic progression are 9 and 93.If the sum of the arithmetic progression is 663,find the number of terms.
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)


Sn = n/2 (a+l)
a= first term
l = last term
663=n/2(9+93)
663*2 = n(102)
663*2/102= n
=13
Number of terms = 13

Answer by stanbon(75887)

The first and the last term of an arithmetic progression are 9 and 93.If the sum of the arithmetic progression is 663,find the number of terms.
---
a = 9
a(n) = 9+(n-1)d = 93
----
(n/2)(9+93) 663
------
n*51 = 663
n = 13
-----
Ans:: n = 13
--------------
Cheers,
Stan H.
-----------------